3. Based on the following costs from a retrospective analysis, what is the 2013 value for the three alternatives using a medical consumer price index (MCPI) inflation rate of 3.5% per year? Year Alternative 1 costs ($) Alternative 2 costs ($) Alternative 3 costs ($) 2011 30,000 15,000 20,000 2013 20,000 25,000 20,000 2010 10,000 15,000 2012 30,000 25,000 20,000
Answer #1

Alternative : 1

10000 X (1.035)3 + 30000 X (1.035)2 + 30000 X (1.035) + 20000 = 94273.93

Alternative : 2

15000 X (1.035)3 + 15000 X (1.035)2 + 25000 X (1.035) + 25000 = 83574.14

Alternative : 3

20000 X (1.035)3 + 20000 X (1.035)2 + 20000 X (1.035) + 20000 = 84298.85

Alternative-2 has the least cost, so it will provide maximum benefit.